OpenStudy (anonymous):

Evaluate of the limit. \[\lim_{x \rightarrow 0}\frac{ \sin3x }{ 5x }\]

OpenStudy (anonymous):

have you tried lhopitals rule

OpenStudy (anonymous):

I didn't learn it yet..

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

trying asking other people while i remember this stuff

OpenStudy (anonymous):

I think it's 5/3

OpenStudy (anonymous):

it is not

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

Oh ok sorry-

OpenStudy (anonymous):

It's 3/5. I read it wrong.

OpenStudy (anonymous):

Sorry I was wrong :/

OpenStudy (anonymous):

I'm not sure how to get 3/5?

OpenStudy (anonymous):

have you learned the maclaurin series

OpenStudy (anonymous):

No, sorry..

OpenStudy (anonymous):

The question is from limit of Trignometric Functions.

OpenStudy (anonymous):

yeah i don't think ill be able to help because i don't remember any of this

OpenStudy (anonymous):

@mathmate

OpenStudy (rational):

you may use below limit \[\large \lim\limits_{t\to 0}~\frac{\sin (t)}{t}=1\]

OpenStudy (rational):

\[\large \lim\limits_{x\to 0}~\frac{\sin (3x)}{5x}\] let \(t=3x\implies x = \frac{t}{3}\) as \(x\to 0\), we have \(t\to 3*0 =0\)

OpenStudy (rational):

the limit becomes : \[\lim\limits_{t\to 0}~\frac{\sin (t)}{5(\frac{t}{3})} = \lim\limits_{t\to 0}~\frac{3\sin (t)}{5t} =\frac{3}{5} \lim\limits_{t\to 0}~\frac{\sin (t)}{t}=\frac{3}{5}\cdot 1=\frac{3}{5}\]

OpenStudy (mathmate):

@esam2 Have you learned the squeeze theorem?

OpenStudy (anonymous):

Yes

OpenStudy (mathmate):

Try example 2 of http://www.analyzemath.com/calculus/limits/squeezing.html

OpenStudy (anonymous):

I think I got it now.

OpenStudy (mathmate):

Great! :)